# Localized Graph-Based Neural Dynamics Models for Terrain Manipulation

### v1

@@ IV-A Implementation Details
Approximately 1,000 robot-terrain interaction trajectories, each uniquely parameterized, were simulated. Throughout these simulations, the downsampled terrain particles were meticulously tracked to facilitate the learning of the dynamics model. Training on this simulation data involved regressing the motion trajectories of the particles using the Mean-Square Error (MSE) criterion, aided by the perfect association of scene particles. Concurrently, the CNN proposer was trained as detailed in Section III-E. The model’s forward simulation steps were set at 0.1 seconds.
IV-B Computational Efficiency and Accuracy
@@
Figure 6: Comparing our method against full-scale GBND (FS) and geometric region proposers with different size (Geo-X). Our method demonstrates significant advantages in speed and GPU memory. $\approx$ 3,000 particles and batch size 256 are used in these experiments, per sample measurements are reported (i.e., divided by 256). Colors and labels are shared by both figures.
@@
The computational performance and accuracy of our method are compared against two baselines. The first baseline utilizes all particles within a large-scale terrain, while the second employs a naïve geometric Region of Interest (RoI), specifically, a cylinder centered at the scoop tip with a predetermined radius. To maintain computational feasibility with the full-terrain baseline, terrains are restricted to 3,000 particles — significantly fewer than what might be encountered by an excavator operating on an large worksite. Figure 6 presents the results on a terrain measuring 50 cm $\times$ 50 cm $\times$ 5 cm, with a particle density of 0.5 particles per cm3. Here, ‘FS’ denotes the full-scale terrain predictor, and ‘Geo-X’ represents the geometric RoI with a cylinder diameter of X cm. Each RoI-based approach is developed using the methodology described in Section III-E alongside our novel node boundary features.
@@
In terms of per-sample processing time and GPU memory utilization, our method matches the speed of the smallest geometric RoI and requires even less GPU memory, while delivering superior accuracy. It is important to note that enhanced RoI prediction significantly improves the accuracy of dynamics predictions. This is because the GBND is prone to inaccuracies, such as predicting minimal movements of static particles. An RoI introduces a strong inductive bias that is absent in full-scale GBND, leading to jitter among distant particles. For smaller values of X, Geo-X exhibits large errors due to the exclusion of moving particles. Conversely, as X increases, the inclusion of numerous static particles leads to performance degradation due to GBND jitter.
Batch Ours Geo-6 Geo-12 Geo-18 FS SAPIEN
@@
TABLE I: Average prediction time (ms) per sample as the batch size varies, showing that 256 is an optimal batch size for our framework. For SAPIEN, 3000 cubes are simulated.
Furthermore, since the planning speed and quality are heavily influenced by the batch size, memory efficiency becomes a critical objective for our dynamics model. Table I illustrates that our method reaches optimal performance at a batch size of $b=256$, achieving an average processing time of less than 0.2 ms per sample. Under these conditions, the planner is capable of rolling out hundreds of trajectories across numerous time steps in seconds. We also compare against the recently updated GPU-based simulator in SAPIEN 3.0 [37] in scenes with 3,000 cube-shaped particles, without a simulated robot. Our method is an order of magnitude faster and can handle larger batch sizes within available GPU memory. More importantly, it also allows fine-tuning to match real-world material behavior.
IV-C Real World Terrain Manipulation
@@
Figure 7: Materials used in real world experiments, U.S. quarter coin on the bottom right for scale. Figure borrowed from [1].
To adapt our model to real-world data, we sampled 100 parametrized robot-terrain interaction trajectories, monitored via an overhead RGB-D camera. Scene point clouds were derived from depth images, with robot proprioception utilized to exclude robot particles while selectively retaining in-tool (scoop) particles by re-sampling within the tool’s convex hull. Due to the inability to track individual particle trajectories across frames, measures such as the weighted sum of chamfer distance [41] and earth-mover’s distance [42] were employed. This fine-tuning process typically improves the dynamics models’ prediction accuracy by approximately 5–10%, compared to the simulation-trained model (Fig. 8).
@@
Figure 8: Fine-tune (FT) the pre-trained (PT) GBND model on real-world datasets significantly improves the model stability and precision in practical scenarios, as evidenced by a reduced median error and a more compact interquartile range.
@@ V CONCLUSION AND LIMITATION
We presented a framework for handling unbounded terrains in graph-based neural dynamics (GBND) models that simultaneously learns a RoI and particle dynamics. Novel node features are key to success of this model. Experiments show that this method achieves orders of magnitude faster prediction and GPU memory usage compared with naïve GBND models or GPU-based physics simulators. It can be trained with both simulated and real data, and when used as a model for planning our method can achieve desired target shapes on real-world excavation platforms.
There remains some limitations of our work that we would like to address in future work. First, our RoI proposer does not accurately model long-range terrain effects (e.g., landslides). Second, our current platform uses an overhead camera to update the terrain, and we would like to extend this to egocentric 3D mapping, in which our model may need to model occluded parts of the terrain. Third, our models assume terrain homogeneity, and future work should address heterogeneous materials including terrains with embedded obstacles. Lastly, the current dynamics model requires full-parameter fine-tuning when moving to unseen materials, one potential solution is to adapt physics-parameter conditioned GBND [44] for online adaptation.

### v2

@@ IV-A Implementation Details
Approximately 1,000 robot-terrain interaction trajectories, each uniquely parameterized, were simulated. Throughout these simulations, the downsampled terrain particles were meticulously tracked to facilitate the learning of the dynamics model. Training on this simulation data involved regressing the motion trajectories of the particles using the Mean-Square Error (MSE) criterion, aided by the perfect association of scene particles. Concurrently, the CNN proposer was trained as detailed in Section III-C. The model’s forward simulation steps were set at 0.1 seconds.
IV-B Particle-based v.s. 2D Heightmap Representations
Figure 5: Rollouts (left to right) from a learned 2D CNN heightmap-based dynamics model compared to L-GBND and ground truth. In the CNN, fine-grained features are lost due to the compressed representation and smoothing bias of CNNs. In contrast, L-GBND preserves volumetric structure and local interactions for 37% lower prediction error on a test dataset.
We compare L-GBND against a baseline that represents terrain using a 2D heightmap and learns dynamics directly in heightmap space. Both models are trained on the same dataset using the same control inputs. For consistency and implementation efficiency, the 2D baseline reuses the architecture of our RoI proposing CNN: a U-Net-style convolutional network that takes as input a stack of past terrain heightmaps and current control input, and autoregressively predicts the future heightmap. To ensure a fair comparison, we project the output of the particle-based model onto a heightmap before computing evaluation metrics. We evaluate prediction accuracy using the $L_{1}$ loss over the heightmap grid. Quantitatively, L-GBND achieves approximately 37% lower prediction error compared to the 2D heightmap model. Fig. 5 presents qualitative results comparing heightmap-based dynamics prediction and ground-truth terrain heightmaps.
This comparison highlights the critical role of representation in deformable terrain dynamics modeling. The 3D particle-based representation used in GBND captures volumetric structure and encodes fine-grained geometry, providing strong inductive biases toward spatial locality, directional interaction patterns, and mass conservation. These properties enable the model to represent contact-rich interactions such as scooping, penetration, and material buildup along tool edges—phenomena that are difficult to express in 2D. In contrast, 2D heightmap representations collapse the terrain into a single surface, discarding sub-surface interactions and occluding contact geometry. Moreover, convolutional architectures tend to produce overly smoothed predictions, which dampens sharp discontinuities and local terrain features. These limitations manifest even more clearly when directly predicting future terrain shape.
IV-C L-GBND’s Efficiency and Effectiveness
@@
Figure 6: Comparing our L-GBND against GBND and geometric region proposers with different size (Geo-X). Our method demonstrates significant advantages in speed and GPU memory. $\approx$ 3,000 particles and batch size 256 are used in these experiments, per sample measurements are reported (i.e., divided by 256). Colors and labels are shared by both figures.
@@
The computational performance and accuracy of our method are compared against two baselines. The first baseline utilizes all particles within a large-scale terrain, while the second employs a naïve geometric Region of Interest (RoI), specifically, a cylinder centered at the scoop tip with a predetermined radius. To maintain computational feasibility with the full-terrain baseline, terrains are restricted to 3,000 particles — significantly fewer than what might be encountered by an excavator operating on an large worksite. Figure 6 presents the results on a terrain measuring 50 cm $\times$ 50 cm $\times$ 5 cm, with a particle density of 0.5 particles per cm3. Here, ‘Geo-X’ represents the geometric RoI with a cylinder diameter of X cm. Each RoI-based approach is developed using the methodology described in Section III-C alongside our novel node boundary features.
@@
In terms of per-sample processing time and GPU memory utilization, our method matches the speed of the smallest geometric RoI and requires even less GPU memory, while delivering superior accuracy. It is important to note that enhanced RoI prediction significantly improves the accuracy of dynamics predictions. This is because the GBND is prone to inaccuracies, such as predicting minimal movements of static particles. An RoI introduces a strong inductive bias that is absent in GBND, leading to jitter among distant particles. For smaller values of X, Geo-X exhibits large errors due to the exclusion of moving particles. Conversely, as X increases, the inclusion of numerous static particles leads to performance degradation due to GBND jitter.
Furthermore, since the planning speed and quality are heavily influenced by the batch size, memory efficiency becomes a critical objective for our dynamics model. Table II illustrates that our method reaches optimal performance at a batch size of $b=256$, achieving an average processing time of less than 0.2 ms per sample. Under these conditions, the planner is capable of rolling out hundreds of trajectories across numerous time steps in seconds. We also compare against the recently updated GPU-based simulator in SAPIEN 3.0 [43] in scenes with 3,000 cube-shaped particles, without a simulated robot. Our method is an order of magnitude faster and can handle larger batch sizes within available GPU memory. As a result of superior efficiency, L-GBND leads to better planning performance in terrain shaping tasks as shown in Table III, where we analyze the scaling of planning performance with respect to number of candidate trajectories being simulated.
Batch L-GBND Geo-6 Geo-12 Geo-18 GBND SAPIEN
@@
TABLE II: Average prediction time (ms) per sample as the batch size varies, showing that 256 is an optimal batch size for our framework. For SAPIEN, 3000 cubes are simulated.
Batch 1 2 4 8 16 32 64 128 256 512 1024
# rollouts 216 407 714 1163 1916 2967 4310 5587 6173 5525 5051
$L_{1}(10^{-3})$ 15.32 14.53 12.68 11.37 10.91 7.22 5.72 3.93 3.42 4.11 4.31
TABLE III: For given time budget (1 sec), more samples being simulated with L-GBND in MPPI planning leads to better planning performance.
IV-D Real World Terrain Manipulation
@@
To evaluate real-world transfer, we fine-tune the pretrained model on 100 real-world interaction trajectories collected via autonomous self-play and overhead RGB-D sensing. Following the procedure described in Section III-E, we compute supervision signals using Chamfer and Earth Mover’s Distance between predicted and observed point clouds. This fine-tuning enables the model to adapt effectively to real-world sensor data and contact dynamics.
@@
Figure 7: Materials used in real world experiments, U.S. quarter coin on the bottom right for scale.
@@
Figure 8: Fine-tuning (FT) the pre-trained (PT) L-GBND model on real-world datasets significantly improves the model stability and precision in practical scenarios, as evidenced by a reduced median error and a more compact interquartile range.
@@ V CONCLUSION AND LIMITATION
We presented L-GBND for handling unbounded terrains in graph-based neural dynamics models that simultaneously learns a RoI and particle dynamics. Novel node features are key to success of this model. Experiments show that this method achieves orders of magnitude faster prediction and GPU memory usage compared with naïve GBND models or GPU-based physics simulators. It can be trained with both simulated and real data, and when used as a model for planning our method can achieve desired target shapes on real-world excavation platforms.
There remains some limitations of our work that we would like to address in future work. First, our RoI proposer does not accurately model long-range terrain effects (e.g., landslides). Second, our current platform uses an overhead camera to update the terrain, and we would like to extend this to egocentric 3D mapping, in which our model may need to model occluded parts of the terrain. Third, our models assume terrain homogeneity, and future work should address heterogeneous materials including terrains with embedded obstacles. Lastly, the current dynamics model requires full-parameter fine-tuning when moving to unseen materials, one potential solution is to adapt physics-parameter conditioning [48] for online adaptation.